OpenStudy (anonymous):

Will fan and medal! @kx2bay (just the top question)

OpenStudy (anonymous):

Q 13 replace x=-2 in each equation what do you get? can you write down your answers for me to verify?

OpenStudy (anonymous):

For the first one y = 4? or do you keep it f(x)=4?

OpenStudy (anonymous):

q 14 - you should be able to answer based on what we covered below, remember where I asked you choose between b and c and you chose which was wrong, well graph (a) here looks almost the same and it is NOT linear. Now I'll get back to q 13

OpenStudy (anonymous):

q 13 a) f(-2) = 4*(-2) = -8 b) p(-2) = 1/2 * (-2) + 2 = -1 + 2 = 1 c) can you try this one... hint the answer will be negative

OpenStudy (anonymous):

Uhmm... I'm doing something wrong here I think g(x) = -2(-2)-1 g(x)= 4 - 1 g(-2) = 3 Can you show me what I'm doing wrong?

OpenStudy (anonymous):

actually you are right and I was wrong, sorry ... your answer is correct

OpenStudy (anonymous):

Oh yay! *happy dance* No problem (:

OpenStudy (anonymous):

happy days

OpenStudy (anonymous):

I'm done! Thank you sooo much for you help!

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

welcome buddy and go to sleep now

OpenStudy (anonymous):

I have to work on Biology now sadly :( Fighting the clock to get as high grades as possible! I am maintaining a 4.05 gpa and it isn't always easy haha

OpenStudy (anonymous):

ohh good luck then, hope u keep the high scores

OpenStudy (anonymous):

Thanks! I do too haha, it's just pesky math!